OpenStudy (anonymous):

intregals of (ln x)^2/x limit e^2 to e^1/2

OpenStudy (solomonzelman):

\[\LARGE \int\limits_{e^{1/2}}^{e^{2}}\ln(x^{2/x})~dx\]correct?

OpenStudy (solomonzelman):

or is it, \[\LARGE \int\limits_{e^{1/2}}^{e^{2}} (\ln x)^{2/x}~dx\]

OpenStudy (solomonzelman):

@isch are you there?

OpenStudy (solomonzelman):

Last attempt, maybe it's \[\LARGE \int\limits_{e^{1/2}}^{e^{2}} \frac{(\ln x)^{2}}{x}~dx\]

OpenStudy (anonymous):

If it's the first or third of those, you're in luck. I don't think the second one has an easily computable result.